OpenStudy (anonymous):

Adult tickets for a play cost $19 and child tickets cost $17. If there were 36 people at a performance and the theatre collected $646 from ticket sales, how many children attended the play?

OpenStudy (anonymous):

let c be the total number of tickets for children let a be the total number of tickets for adults The equations are: c + a = 36 19a + 17c = 646 Solving for c: a = 36 - c 17c + 19(36 - c) = 646 17c - 19c + 684 = 646 -2c = -38 c = 19

OpenStudy (anonymous):

Thanks!